# Superficial Safety Alignment Hypothesis

### v1

@@ 4 Less is More for Safety Alignment
Based on the Superficial Safety Alignment Hypothesis (SSAH), we posit that safety alignment only needs to teach the model the correct reasoning direction - either fulfilling or refusing a request - and to equip it with a standard refusal mechanism. This leads to the insight that safety alignment can be achieved using only a small subset of critical computing units, as the task can be interpreted as a binary classification combined with a multi-selection task.
@@ 4.1 Identifying Safety-Critical Computing Units
To verify this corollary, we designed experiments to determine the minimally essential subset of computing units in a large language model that is critical in establishing a safety guardrail. Following this line of reasoning, we hypothesize that specific attributes of LLMs can be explicitly linked to certain computing units within the model. Our experiments are designed as follows:
Definition of Attribute Groups. This paper categorizes the attributes of LLMs into two main properties: utility and safety. Following the above hypothesis, we first exclusively link safety or utility attributes to specific computing units. We also speculate that some units may contribute to both attributes simultaneously. Moreover, considering that many components in LLMs are redundant, we also hypothesize that certain computing units do not correlate with any attribute. Therefore, we divide the computing units of LLMs into four groups: Exclusive Safety Units (ESU) and Exclusive Utility Units (EUU), which are linked exclusively to either safety and utility, respectively; Complex Units (CU), which concurrently contribute to both safety and utility attributes; and Redundant Units (RU), which are not associated with any attribute.
Table 1: Pruning results of Llama2-7B-Chat and Llama3-8B-Instruct across safety and utility benchmarks. Breakdown of model performance after pruning various categories of computing units, including ESU, EUU, and RU, demonstrating their respective contributions to safety and utility attributes. The proportion of each attribute group in the model is determined based on the degradation in utility and safety. Additional evaluation details are provided in Appendix B.1 and B.2.
@@
Here, $N$ refers to the number of calibration samples, $\mathbf{W}^{B}_{:,j}$ refers to the $j$-th column of the weight matrix $B$, and $X^{B}$ represents the input to $B$. Based on this score, we plan to prune the input channels of $B$ and the output neurons of $A$, as channels or neurons with low activation variance across the target dataset are considered less important. More details can be found in the Appendix B.3.
In this way, we calculate the importance score for each individual neuron or channel, denoted as $\mathbf{I_{U}}$ for the utility attribute and $\mathbf{I_{S}}$ for the safety attribute. Initially, we prune the computing units with the smallest $\mathbf{I_{U}+I_{S}}$ values to identify redundant units. Subsequently, we prune units with the largest and smallest $\mathbf{I_{S}-I_{U}}$ values to identify exclusive utility and safety units, respectively. The remaining computing units are categorized as complex units. We experiment with various pruning ratios and evaluate the resulting safety and utility performance, selecting the optimal pruning ratio with minimal performance degradation in the corresponding attribute. This systematic pruning process enables us to accurately derive the roles of different units and validate our hypothesis regarding safety and utility-critical computing units.
Ablation Study Results. The experiment results are described in Tab. 1. We discovered that for a safety-aligned model, the computing units that are exclusively responsible for the safety attribute account for only about 1.3 - 1.4% of the total units. Although the complex units make up a larger portion of the model, their primary role is to support the general knowledge required for both safety and utility tasks. Based on these findings, we have partially validated our hypothesis that safety alignment is relying on (not constructed by) a subset of safety-critical computing units.
@@ 4.2 Why is Safety Brittle?
Figure 5: Attribute transfer analysis for the downstream task (Dolly Dataset) fine-tuning on Llama2-7B-Chat. More than half of the ESU transferred to CU, while part of the CU transferred to EUU. Although a significant portion of RU transferred to CU, this mainly contributes to utility due to the objective of the fine-tuning task. Overall, the computing units that originally contributed to safety decreased (Transfer portions less than 0.1% are excluded from this figure.)
Attribute Transfer Analysis in the Fine-Tuning Process. Previous research has shown that adapting safety-aligned LLMs to new tasks can often hurt their safety performance. Therefore, it is crucial to examine how and how much the attributes of individual computing units change during this process. To investigate this, we designed an experiment where a safety-aligned LLM is fine-tuned on a downstream dataset, and we track the changes in the attributes of its computing units over time. The transfer statistics are summarized in Fig. 5, revealing several important patterns: First, for a safety-aligned LLM, regardless of whether it has been fine-tuned on a different task, the majority of computing units are classified as complex units. This suggests that safety and utility attributes are deeply intertwined within the aligned model. Second, during the fine-tuning process, a significant number of exclusive safety units and complex units are converted into exclusive utility units. This transformation indicates that fine-tuning for utility tends to shift the function of computing units away from safety, compromising the model’s ability to maintain its safety guardrails.
Based on these observations, we draw the following insight: During the task adaption of LLMs, the model often obtains the expected attribute (e.g., utility) by converting computing units that originally contributed to the other attribute (e.g., safety). This means that enhancing utility performance in a different task often comes at the expense of the safety performance.
Table 2: Safety performance of Llama2-7B-Chat and Llama3-8B-Instruct under Fine-Tuning attacks (Alpaca and Dolly) across various benchmarks and judge methods. We compare the measures of the initial models, fine-tuned models, and our strategies. Specifically, our strategy includes two settings: Setting (i): freeze all ESU and the top 6% of CU, and Setting (ii): freeze all ESU and all CU. Both strategies demonstrate significant mitigation of safety performance degradation. Bold indicates the best results, while the underlined mark the second-best results. Note that we doubled the training epochs for our method to ensure a fair comparison, resulting in identical or lower final training loss compared to the fine-tuned models. Additionally, due to computational limitations, we froze the first 12 transformer blocks of LLaMA3, although similar trends are observed. We also fine-tuned the model with Setting (ii) for one epoch prior to applying Setting (i). Further details are available in Appendix B.5.
@@
Freezing Safety-Critical Components to Preserve Safety Guardrails. Given the above insight, we propose freezing the identified safety-critical components during the fine-tuning process to prevent the unwanted attribute transfer of these units and thereby preserve safety performance. To test this hypothesis, we conduct experiments across different language models, where we freeze the safety-critical components identified through our aforementioned ablation study in the fine-tuning process. The experiment results are described in Tab. 2, and we have the following observations: Observation 1) Brittleness of current safety mechanisms. Adapting a fully aligned model to new tasks significantly increases the attack success rate or harmful output scores across various attack benchmarks and evaluation methods. This clearly demonstrates the brittleness of current safety mechanisms in LLMs. Observation 2) Effectiveness of freezing safety-critical components By freezing the Exclusive Safety Units (ESU) and the top 6% of Complex Units (CU), we can significantly reduce the degradation of safety guardrails during the adaptation process across different models, tasks, benchmarks, and evaluation methods. Freezing all CU components further improves performance (The ablation experiments with varying proportions of frozen CUs are detailed in Appendix C.1). Observation 3) Fragility of safety in newer models. We found that the safety guardrails in LLaMA 3 are more fragile than in LLaMA 2. We speculate it because LLaMA 3 attempts to analyze the true intentions behind harmful requests, which can lead to more errors. This suggests that the community should carefully consider how models should respond to harmful requests, whether to refuse outright or engage in further interactions. Observation 4) Additionally, we conduct an analysis of the attribute transfer in this new setting and observe that freezing the safety-critical components successfully mitigates the conversion of safety units into utility units.
Table 3: Safety performance of Llama2-7B-Chat under Fine-Tuning attacks (Alpaca and Dolly) across various parameter-efficient fine-tuning setups. Results with are from Qi et al. (2023).
Comparing with Parameter-Efficient Fine-Tuning (PEFT) Approaches. To ensure that the preservation of safety in our approach is not merely due to not updating the model itself, we also examine how safety guardrails degrade during the parameter-efficient fine-tuning of LLaMA2-7B. Specifically, we tested three PEFT methods: LoRA, LLaMA-Adapter, and Prefix Tuning. The experimental results are detailed in Table 2, and we found that these methods led to worse degradation of safety compared to full-model fine-tuning and even further so when compared to our approach. This indicates that the effectiveness of preserving safety in our method primarily stems from successfully identifying safety-critical components rather than simply freezing the model.
4.3 Free Lunch: Repurposing RUs as Alignment Budget Even Help Reduce Tax
Up to this point, we have successfully identified safety-critical components and preserved the model’s safety properties while adapting LLMs to new tasks. From another perspective, an interesting question arises: Can we assign safety or utility attributes directly to certain computing units, respectively? If we could, this would mean that we can manage the attributes of LLMs in a fine-grained, controlled way. Building on this, the next straightforward question is whether we can convert previously redundant units (units that do not contribute to the utility performance of a pre-trained model) into safety-critical units that strengthen the safety guardrails of a safety-aligned model.
Importantly, we do not need to convert too many computing units to achieve safety alignment. Previous research has found that at least 20% of the parameters in pretrained LLMs are redundant (Li et al., 2024; Ma et al., 2023; An et al., 2024). This special observation motivates us to consider the following: with such a large percentage of parameters in pre-trained LLMs available as an alignment budget, can we design an alignment method that reduces the alignment tax?
@@ 4.3 Free Lunch: Repurposing RUs as Alignment Budget Even Help Reduce Tax
Figure 6: Attribute transfer analysis for alignment on Llama2-7B and Llama2-7B-Chat. A significant portion of computing units that originally contributed solely to utility is flipped to play a more comprehensive role after alignment. (Transfer portions less than 0.1% are excluded.)
Attribute Transfer Analysis in the Alignment Process. Before testing the above hypothesis, we conducted an attribute transfer analysis during the alignment process to explore the underlying reasons for the alignment tax. For the pre-trained LLM, we simply categorized the attribute groups into Utility Units and Redundant Units, as the model has not yet undergone safety alignment. This is different from the categorization we used for aligned LLMs. The transfer statistics are summarized in Fig. 6, and they reveal the following key pattern: a large percentage of units that originally contributed to utility in the pre-trained model are transferred to CU and ESU in the aligned LLM. In contrast, the originally redundant units remain largely unused after the alignment process.
With this observation, we are strongly motivated to verify the above question. Specifically, we use the pruning method described in Sec. 4.1 to identify the redundant units in Llama-7B, as there is no officially aligned model for it. Once we identify the redundant units, we freeze the updates for the rest of the model’s parameters and perform fine-tuning only on these redundant units. To reduce the complexity of the experiment, we focus directly on the general alignment process instead of safety alignment, since the results should hold for the latter as a subset. The experiment results are shown in Table 4, and we successfully implemented alignment with only 20% parameter updates and without incurring alignment tax, especially highlighting the mathematical performance. These findings have significant implications for the scalability and efficiency of safety alignment in future LLMs.
@@
5 Discussion, Limitation, and Conclusion
Discussion. While our SSAH offers valuable insights into adversarial scenarios, such as jailbreak attacks, we do not propose a specific solution to address these issues in this work. If these issues could be resolved within the framework of our theory, the term “Superficial” in “Superficial Safety Alignment Hypothesis” may no longer be necessary. Interestingly, recent research provides some supporting evidence in this direction (Qi et al., 2024). However, it is also highly likely that advanced attacks may not be fully mitigated by relying solely on the model’s internal mechanisms. A systematic, multi-layered approach, extending beyond the model itself, may be required to effectively defend against sophisticated adversarial threats.
Limitation. When reallocating redundant units for safety purposes, we only explored the impact of the alignment method SFT. Due to resource limitations, we have not yet tested this approach on other alignment methods like PPO or DPO.
@@ 5 Discussion, Limitation, and Conclusion
Conclusion. This paper distinguishes safety alignment from the general alignment in LLMs and addresses the three key questions: How does safety alignment affect model behavior? Why are safety mechanisms brittle? and How to mitigate the safety alignment tax? By answering these questions, we were able to demonstrate that safety alignment can be a straightforward process, rather than a myth.
@@ A.2 Model Configuration and Training Details
Figure 7: Probing reasoning direction on the Hex dataset with Llama2-7B, Llama3-8B, and Llama3.1-8B using cosine distance. Models were fine-tuned to ensure that aligned versions possess both general instruction-following abilities and safety guardrails, while unaligned models only have instruction-following capabilities. More results and details can be found in Appendix A.2.
Thus, for each LLaMA variant (LLaMA2, LLaMA3, LLaMA3.1), we fine-tuned two separate models using Supervised Fine-Tuning (SFT):
  1. (1)
    A General Instruction-Following Model that is trained to follow human instructions but without any explicit safety mechanisms. This model helps us evaluate how a model with instruction-following capabilities but without safety guardrails reacts to malicious queries.
  2. (2)
    A Safety-Aligned Model that incorporates both general instruction-following capabilities and explicit safety mechanisms, allowing us to examine how safety alignment influences the model’s reasoning direction when responding to unsafe inputs.
By comparing these two categories of models when exposed to different types of malicious queries, we can better understand how safety alignment reshapes the internal decision-making process of large language models.
@@
We follow the alignment method outlined in the Zhou et al. (2024), which uses Supervised Fine-Tuning (SFT). For the general instruction-following models, we employed the LIMA dataset, which includes over 1000 instruction-following examples. However, we removed 13 safety-related examples to avoid conflating safety concerns with general instruction-following abilities. The filtering process was assisted by GPT-4, following a set of instructions specifically designed to identify and exclude safety-related tasks. For the safety-aligned models, we used the Alert dataset, which contains a variety of safety-critical instructions to teach the model how to respond safely to malicious queries.
For all models, we followed a consistent training configuration across the LLaMA2, LLaMA3, and LLaMA3.1 versions to ensure comparable results. The general instruction-following models were trained on the LIMA dataset (with safety-related data removed). The fine-tuning was performed using the following key parameters:
  • Batch size: We set the batch size per device to 4, with gradient accumulation steps of 6 on 3 NVIDIA A6000 GPU, which gave us an effective batch size of 72.
  • Learning rate: The learning rate was set to 1.0e-5.
  • Epochs: The fine-tuning was conducted for 15 epochs.
  • Precision: BF16 precision was used to optimize memory usage.
  • Optimization: AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of 0.1.
  • Learning rate scheduler: We employed a linear scheduler with no warm-up steps.
  • Seed: A random seed of 42 was used for reproducibility.
For the safety-aligned models, we fine-tuned the instruction-following models (trained on the LIMA dataset) further using the Alert dataset. The key fine-tuning parameters for this stage were as follows:
  • Model initialization: We initialized the model from the previously fine-tuned general instruction-following model.
  • Batch size: As with the general instruction model, we used a batch size of 4 with gradient accumulation steps of 6 on 3 NVIDIA A6000 GPU.
  • Learning rate: The learning rate was set to 1.0e-5.
  • Epochs: We fine-tuned the safety-aligned models for 9 epochs.
  • Precision: BF16 precision was used to optimize memory usage.
  • Optimization: The same AdamW optimizer configuration was applied.
By fine-tuning both general instruction-following and safety-aligned models with these configurations, we create a controlled environment for probing the reasoning direction in response to various queries. This enables us to systematically compare the behaviors of different model types under the same conditions and assess the impact of safety alignment at each generation step.

### v2

@@ 4 Less is More for Safety Alignment
While the full validation of SSAH remains challenging due to the high complexity of the output sampling space, we are able to empirically validate a key corollary. Specifically, if SSAH holds, then safety alignment can be achieved with only a small subset of critical computing units, as the task can be interpreted as an implicit and simple safety-related binary classification problem.
@@ 4.1 Identifying Safety-Critical Computing Units
To verify this corollary, we designed experiments to determine the minimally essential subset of computing units in LLMs that is critical in establishing a safety guardrail. Following this, we hypothesize that specific attributes of LLMs can be explicitly linked to certain computing units within the model. Our experiments are designed as follows:
Definition of Attribute Groups. This paper considers two attributes of LLMs, utility and safety. We first exclusively link safety or utility attributes to specific computing units. We also speculate some units may contribute to both attributes simultaneously. Moreover, we hypothesize that certain computing units are redundant and do not associate with any attribute. Therefore, we categorize the computing units of LLMs into four groups: Safety Critical Units (SCU) and Utility Critical Units (UCU), corresponding exclusively to either safety and utility, respectively; Complex Units (CU), contributing to both safety and utility; and Redundant Units (RU), not associated with any attribute.
Table 1: Pruning results of Llama2-7B-Chat and Llama3-8B-Instruct across safety and utility benchmarks. Breakdown of model performance after pruning various categories of computing units, including SCU, UCU, and RU, demonstrating their respective contributions to safety and utility attributes. The proportion of each attribute group in the model is determined based on the degradation in utility and safety. Additional evaluation details are provided in Appendix B.1 and B.2.
@@
Here, $N$ refers to the number of calibration samples, $\mathbf{W}^{B}_{:,j}$ refers to the $j$-th column of $B$, and $X^{B}$ represents the input to $B$. With this score, we plan to prune the input channels of $B$ and the output neurons of $A$, as channels or neurons with low activation variance across the target dataset are considered less important - details in Appendix B.3.
In this way, we calculate the importance score for each individual neuron or channel, denoted as $\mathbf{I_{U}}$ for the utility attribute and $\mathbf{I_{S}}$ for the safety attribute. Initially, we prune the computing units with the smallest $\mathbf{I_{U}+I_{S}}$ values to identify redundant units. Subsequently, we prune units with the largest and smallest $\mathbf{I_{S}-I_{U}}$ values to identify utility and safety critical units, respectively. The remaining computing units are categorized as complex units. We experiment with various pruning ratios (starts with higher ratios and systematically reduces them) and evaluate the resulting safety and utility performance, selecting the optimal pruning ratio with minimal performance degradation in the corresponding attribute. This systematic pruning process enables us to validate our hypothesis and derive the roles of different units accurately.
Ablation Study Results. The experiment results are described in Tab. 1. We discovered that for a safety-aligned model, the computing units that are exclusively responsible for the safety attribute account for only about 1.3 - 1.4% of the total units. Although the complex units make up a larger portion of the model, their primary role is to support the general knowledge required for both safety and utility tasks. Based on these findings, we have partially validated our hypothesis that safety alignment is relying on (not constructed by) a subset of safety-critical computing units. We also provide more experiment results in Appendix C.
@@ 4.2 Why is Safety Brittle?
Figure 5: Attribute transfer analysis for the downstream task (Dolly Dataset) finetuning on Llama2-7B-Chat. More than half of the SCU transferred to CU, while part of the CU transferred to UCU. Although a significant portion of RU transferred to CU, this mainly contributes to utility due to the objective of the finetuning task. Overall, the computing units that originally contributed to safety decreased (Transfer portions less than 0.1% are excluded from this figure.)
Attribute Transfer Analysis in the Fine-Tuning Process. Previous research has shown that adapting safety-aligned LLMs to new tasks can often hurt their safety performance. Therefore, it is crucial to examine how and how much the attributes of individual computing units change during this process. To investigate this, we designed an experiment where a safety-aligned LLM is fine-tuned on a downstream dataset, and we track the changes in the attributes of its computing units over time. The transfer statistics are summarized in Fig. 5, exhibiting important observations: First, for a safety-aligned LLM, regardless of whether it has been finetuned on a different task, the majority of computing units are classified as complex units. This suggests that safety and utility attributes are deeply intertwined within the aligned model. Second, during the finetuning process, a significant number of safety critical units and complex units are converted into utility critical units. This transformation indicates that finetuning for utility tends to shift the function of computing units away from safety, compromising the model’s ability to maintain its safety guardrails.
Based on these observations, we draw the following insight: During the task adaption of LLMs, the model often obtains the expected utility by converting computing units that originally contributed to the other attribute safety. This means that enhancing utility performance in a different task comes at the expense of the safety performance.
Table 2: Safety performance of Llama2-7B-Chat and Llama3-8B-Instruct under Fine-Tuning attacks (Alpaca and Dolly) across various benchmarks and judge methods. We compare the measures of the initial models, finetuned models, and our strategies including two settings: setting (i): freezing all SCU and the top 6% of CU, and setting (ii): freezing all SCU and all CU. Both strategies demonstrate significant mitigation of safety performance degradation. Bold indicates the best results, while the underlined mark the second-best results. Note that we doubled the training epochs for our method to ensure a fair comparison, resulting in identical or lower final training loss compared to the finetuned models. Additionally, due to computational limitations, we froze the first 12 transformer blocks of LLaMA3, although similar trends are observed. We also finetuned the model with Setting (ii) for one epoch prior to applying Setting (i). Further details are in Appendix B.5.
@@
Table 3: Utility evaluation across three model variants: base model, fully finetuned model, and finetuned model with our policy. Results are reported on 10 downstream tasks for both LLaMA2-7B-Chat (finetuned on Alpaca) and LLaMA3-8B-Instruct (finetuned on Dolly). Our freezing method does not hurt model utility performance.
Model Variant ARC-C ARC-E BoolQ HellaSwag OpenBookQA PIQA RTE Winogrande GSM8K MMLU
LLaMA2-7B-Chat
Base 44.28 73.91 79.79 75.50 43.60 77.26 69.68 66.38 22.97 46.35
Alpaca 47.44 76.52 81.47 74.68 44.00 79.11 69.68 67.72 19.79 46.82
Alpaca + Freeze (Ours) 47.87 77.19 80.37 74.74 44.40 78.78 68.95 67.56 19.64 47.94
LLaMA3-8B-Instruct
Base 56.74 81.65 83.09 75.85 43.00 78.62 67.51 72.06 75.59 63.92
Dolly 55.63 83.25 80.86 77.22 44.40 80.25 66.43 73.64 66.94 63.43
Dolly + Freeze (Ours) 57.85 83.42 80.34 77.63 44.60 80.36 67.51 73.64 68.39 63.23
@@
Figure 6: Attribute transfer analysis for alignment on Llama2-7B and Llama2-7B-Chat. A significant portion of computing units that originally contributed solely to utility is flipped to play a more comprehensive role after alignment.
Freezing Safety-Critical Components to Preserve Safety Guardrails. Given the above insight, we propose freezing the identified safety-critical components during the finetuning process to prevent the unwanted attribute transfer of these units and thereby preserve safety performance. To test this hypothesis, we conduct experiments across different language models, where we freeze the safety-critical components identified through our aforementioned ablation study in the finetuning process. The experiment results are described in Tab. 2 & 3, and we have the following observations: Observation 1) Brittleness of current safety mechanisms. Adapting a fully aligned model to new tasks significantly increases the attack success rate or harmful output scores across various attack benchmarks and evaluation methods. This clearly demonstrates the brittleness of current safety mechanisms in LLMs. Observation 2) Effectiveness of freezing safety-critical components. By freezing the Safety Critical Units (SCU) and the top 6% of Complex Units (CU), we can significantly reduce the degradation of safety guardrails during the adaptation process across different models, tasks, benchmarks, and evaluation methods. Freezing all CU components further improves performance (The ablation experiments with varying proportions of frozen CUs are detailed in Appendix D.1). Observation 3) Fragility of safety in newer models. We found that the safety guardrails in LLaMA 3 are more fragile than in LLaMA 2 (xie2025sorry). We speculate this because LLaMA 3 attempts to analyze the true intentions behind harmful requests, which can lead to more errors. This suggests that the community should carefully consider how models should respond to harmful requests, whether to refuse outright or engage in further interactions. Observation 4) Additionally, we conduct an analysis of the attribute transfer in this new setting and observe that freezing the safety-critical components mitigates the conversion of safety units into utility units. We provide more experiment results for different model families and downstream datasets in Appendix C.
@@
Comparing with Parameter-Efficient Fine-Tuning (PEFT) Approaches. To verify that the preservation of safety in our approach is not merely due to not updating the model, we also examine how safety performance degrades during the parameter-efficient finetuning of LLaMA2-7B. We tested three PEFT methods: LoRA, LLaMA-Adapter, and Prefix Tuning. The results are detailed in Table 5, and we found that these methods led to worse degradation of safety compared to full-model finetuning and even further when compared to our approach. This indicates that the effectiveness of preserving safety in our method primarily stems from successfully identifying safety-critical components rather than simply freezing the model.
Table 5: Safety performance of Llama2-7B-Chat under Finetuning attacks (Alpaca and Dolly) across various parameter-efficient fine-tuning setups. Results with are excerpted from Qi et al. (2023).
4.3 Repurposing Redundancy as Alignment Budget
Up to this point, we have empirically validated a key corollary of SSAH: less is more in terms of parameters for safety alignment. This naturally leads to a deeper question: Can safety attributes be directly assigned to specific computing units? If this is possible, it opens the door to fine-grained, controllable safety management in LLMs. Building on this idea, a compelling follow-up question would be whether we can convert previously redundant units into safety-critical units that strengthen the safety of LLMs or not.
Previous research has found that at least 20% of the parameters in pretrained LLMs are redundant (Li et al., 2024; Ma et al., 2023; An et al., 2024). This special observation motivates us to consider the following: with such a large percentage of parameters in pre-trained LLMs available as an alignment budget, can we design an alignment method that reduces the alignment tax?
Attribute Transfer Analysis in the Alignment Process. Before testing the above hypothesis, we conducted an attribute transfer analysis during the alignment process to explore the underlying reasons for the alignment tax. For the pretrained LLM, we simply categorize the attribute groups into Utility Units and Redundant Units, as the model has not yet undergone safety alignment. This is different from the categorization we used for aligned LLMs. The transfer statistics are summarized in Fig. 6, and they reveal the following key pattern: a large percentage of units that originally contributed to utility in the pre-trained model are transferred to CU and SCU in the aligned LLM, while the originally redundant units remain largely unused.
With this observation, we are strongly motivated to verify the question above by employing pruning method described in Sec. 4.1 to identify the redundant units in Llama-7B, as there is no officially aligned model for it. Once we identify the redundant units, we freeze the updates for the rest of the model’s parameters and perform finetuning only on these redundant units. To reduce the complexity, we focus directly on the general alignment process instead of safety alignment, since the results should hold for the latter as a subset. The experiment results are shown in Table 4, and we successfully implement the same level alignment with only 20% parameter updates and without incurring alignment tax, especially highlighting the mathematical performance. These findings provide direct evidence for the corollary—less is more in terms of parameters for safety alignment, thereby offering indirect yet compelling support for the validity of SSAH.
5 Broader Impact, Limitations, and Conclusion
Extension to Jailbreak Attacks. SSAH was introduced to explain how existing safety alignment techniques influence model behavior and, from another perspective, to highlight the key limitation—namely, the inability to provide safety guardrails at each generation step. This superficiality makes aligned models struggle against various jailbreak attacks (Qi et al., 2024; li2025superficial). As discussed in Sec. 3, we have outlined a theoretical direction for achieving full safety alignment. Finally, we would extend SSAH to a non-superficial framework:
Given an unsafe model that is capable of fulfilling users’ malicious requests, safety alignment teaches the model to choose and maintain the correct reasoning direction at each generation step, along with simple refusal mechanisms. That is, the model is able to continuously re-evaluate and re-select its reasoning direction—safe or unsafe—throughout the generation process
Limitation and Future Work. Our investigation into reallocating redundant units for safety purposes is limited to studying the effects of SFT on limited alignment datasets. While this is sufficient to offer indirect support for SSAH, future research should expand this line of inquiry by exploring broader datasets and alternative alignment strategies to further validate and generalize our findings.
@@ 5 Broader Impact, Limitations, and Conclusion
Conclusion. This paper distinguishes safety alignment from general alignment and addresses the three key questions: How does safety alignment affect model behavior? Why are safety mechanisms brittle? and How to mitigate the side effects?. By answering these questions, safety alignment is demonstrated as an implicit binary classification task.
@@ A.2 Model Configuration and Training Details
Figure 7: Probing reasoning direction on the Hex dataset with Llama2-7B, Llama3-8B, and Llama3.1-8B using cosine distance. Models were fine-tuned to ensure that aligned versions possess both general instruction-following abilities and safety guardrails, while unaligned models only have instruction-following capabilities.
Thus, for each LLaMA variant (LLaMA2, LLaMA3, LLaMA3.1), we fine-tuned two separate models using Supervised Fine-Tuning: (A) General Instruction-Following Model: a model that is trained to follow human instructions but without any explicit safety mechanisms. This model helps us evaluate how a model with instruction-following capabilities but without safety guardrails reacts to malicious queries. (B) Safety-Aligned Model: a model that incorporates both general instruction-following capabilities and explicit safety mechanisms, allowing us to examine how safety alignment influences the model’s reasoning direction when responding to unsafe inputs. By comparing these two categories of models when exposed to different types of malicious queries, we can better understand how safety alignment reshapes the internal decision-making process of large language models.
@@
We follow the alignment method outlined in the Zhou et al. (2024), which employes Supervised Fine-Tuning (SFT). For general instruction-following models, we used the LIMA dataset, which consists of over 1000 instruction-following examples. However, to ensure that safety-related concerns did not interfere with general instruction-following abilities, we removed 13 safety-related examples. The filtering process was assisted by GPT-4, following a set of instructions specifically designed to identify and exclude safety-related tasks. For safety-aligned models, we utilized the Alert dataset, which contains a set of safety-critical instructions designed to teach models how to respond appropriately to malicious queries.
To maintain consistency across experiments, we applied the same training configuration for all models, including those based on LLaMA2, LLaMA3, and LLaMA3.1. The general instruction-following models were fine-tuned on the modified LIMA dataset (excluding safety-related data), using the following training parameters: a batch size of 4 per device, with gradient accumulation steps of 6 on three NVIDIA A6000 GPUs, resulting in an effective batch size of 72; a learning rate of $1.0\times 10^{-5}$; a total of 15 training epochs; and BF16 precision to optimize memory usage. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of 0.1. The learning rate was scheduled linearly with no warm-up steps, and a random seed of 42 was set to ensure reproducibility. For safety-aligned models, we further fine-tuned the instruction-following models trained on the LIMA dataset using the Alert dataset. The fine-tuning process followed a similar setup, with the model initialized from the previously trained general instruction-following model. The training parameters remained largely consistent, including a batch size of 4 with gradient accumulation steps of 6 on three NVIDIA A6000 GPUs, a learning rate of $1.0\times 10^{-5}$, and BF16 precision. However, the number of training epochs was reduced to 9 to prevent overfitting. The AdamW optimizer was used with the same configuration as in the previous stage.
By fine-tuning both general instruction-following models and safety-aligned models using these controlled conditions, we establish a controlled environment for evaluating examine their reasoning direction when responding to different queries. This approach allows for an indirect comparison of model behaviors under identical conditions, providing insight into how safety alignment influences decision-making at each generation step.